An educational professional wants to determine the number of times students procrastinate on doing homework each week. They designed a chart for students to log their data. Find number of students that will be needed​ (the sample​ size) in order to be 90 percent confident that the sample mean will not differ from the population mean by more than 2 times.​ (The professional estimated the standard deviation to be 3.25 which was found from an earlier​ study.)
Answer #1

Solution :

Given that,

Population standard deviation = \sigma = 3.25

Margin of error = E = 2

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

sample size = n = (Z\alpha/2* \sigma / E) 2

n = (1.645 * 3.25/ 2)2

n = 7.15

n = 8

Sample size = 8